OpenStudy (anonymous):

A sample of 100 light bulbs is tested for failures. Of the 100 light bulbs, 10 of them don’t turn on. The population proportion is .15. Solve for the sample proportion and the standard deviation.

OpenStudy (perl):

I will help you get started. The sample proportion is given as $$ \Large \hat p = \frac{10}{100}$$ The mean of the distribution of sample proportions of size n is : $$ \Large \mu_\hat p = p $$ The standard deviation of the distribution of sample proportions of size n is: $$ \Large \sigma_\hat p =\sqrt{\frac{p(1-p)}{n}} $$

OpenStudy (perl):

note that $$\Large \hat p , p , \mu_{ \hat p } $$ stand for sample proportion, population proportion, population mean of the distribution of sample proportions.

OpenStudy (anonymous):

So what do i do first